I need help with a word problem.King Kong and Godzilla are destroying cities.The ratio of cities destroyed by King Kong to those destroyed by Godzilla is 2:5.If the two together destroy a total of 35 cities,find how many cities each destroyed.Include a check.Thank you for your help.
Answer by checkley79(3341)

2+5=7 THE TOTAL NUMBER OF EVENTS.
35/7=5 ITEMS PER EVENT.
2*5=10 DESTRUCTIONS FOR KING KONG.
5*5=25 DESTRUCTIONS FOR GODZILLA.
PROOF:
10+25=35
2/5=10/25
2/5=2/5